# RLCAD: Reinforcement Learning Training Gym for Revolution Involved CAD Command Sequence Generation

### v1

@@ Abstract
In this paper, we present a reinforcement learning (RL) training environment (gym) built on a CAD geometric engine. Given an input boundary representation (B-Rep) geometry, the policy network in the RL algorithm generates an action. This action, along with previously generated actions, is processed within the gym to produce the corresponding CAD geometry, which is then fed back into the policy network. The rewards, determined by the difference between the generated and target geometries within the gym, are used to update the RL network. Our method supports operations beyond sketches, Boolean, and extrusion, including revolution operations. With this training gym, we achieve state-of-the-art (SOTA) quality in generating command sequences from B-Rep geometries. In addition, our method can significantly improve the efficiency of command sequence generation by a factor of $39X$ compared with the previous training gym.
@@ 1 Introduction
Figure 1: Demonstration of various CAD models generated by RLCAD. Including of revolution operation enables the generation of more complex geometric shapes, such as the battery, nuts, chess pieces, etc.
In our daily lives and industrial production, objects ranging from small items like cups to large structures like airplanes are designed using a series of 2D and 3D modeling operations in computer-aided design (CAD) software. The process of recording these parametric operations results in a command sequence that encapsulates domain expert knowledge by precisely defining a sequence of modeling steps to generate CAD geometries. However, obtaining the command sequence directly is often not feasible. Instead, we should generate the corresponding command sequence using other modalities, such as boundary representation, and point clouds.
@@
Traditional CAD geometry generation algorithms rely on CAD geometric engines, which provide essential functionalities such as validity checks and constraint solving – capabilities that supervised learning alone cannot achieve. Meanwhile, SOTA large language models such as DeepSeek r1 [19] and OpenAI o1 [20] employ RL techniques to enhance text generation quality and model reasoning capabilities. In bionic robotics, RL-based simulation training has also yielded impressive results in robot motion control [21] and manipulation [22].
Inspired by these advancements, we introduce a CAD RL training gym based on a CAD geometric engine (RLCAD). In this framework, the policy network of the RL framework generates a command sequence, which is executed in the gym to produce the corresponding CAD geometry. This geometry is then fed back into the policy network to calculate rewards and update the network.
Based on this training gym, we propose a RL-based CAD geometry generation algorithm that converts B-Rep models into command sequences. Our approach supports operations beyond sketches, Boolean, and extrusion, including revolution operations. As shown in Figure 1, including revolution operation enables the generation of more complex geometric shapes, such as the battery, nuts, chess pieces, etc. We summarize our contributions as follows:
@@
    We present a CAD RL training gym based on a CAD geometric engine, bridging RL algorithms with CAD geometric engines. We provide a high-level Python interface wrapper and support for parallel environment sampling, enabling the creation and interaction with multiple CAD environments.
@@
    We present a RL-based CAD generation algorithm that converts B-Rep models into command sequence. We design a multi-modal policy network integrating revolution operations with cross-modal feature extractors and hybrid-reward functions.
  • 3.
    We integrate RL algorithm library Stable Baseline3 [23] and the OpenAI Gymnasium interface [24] with RLCAD, making it easier for researchers and developers to implement, test, and benchmark RL algorithms.
@@ 2.1 CAD Datasets
Current parametric CAD models typically are stored as B-Rep  [25] and/or command sequence [26]. B-Rep includes the analytical representation of the surface and curves of a 3D shape, while command sequences record the historical construction process of the shape. B-Rep models consist of trimmed parametric surfaces along with topological information describing their adjacency relationships.
@@ 2.2 CAD Model Generation
BrepGen [7] combines vector variational autoencoders and latent space diffusion models for the autoregressive generation of B-Rep models. CADSIGNet [8] employs layer-wise cross-attention between point clouds and command sequences to learn implicit representations of CAD models, enabling end-to-end point cloud to command sequence generation. HNC-CAD [14], SkexGen [15], and DeepCAD [16] quantize and encode CAD command sequences, generating corresponding latent space features and training through an encoder-decoder autoregressive model. SfmCAD [13] uses voxels as input and proposes a differentiable sketch and sweeping path modeling operations to reconstruct the CAD model in a self-supervised manner.
@@ 2.3 B-Rep to Command Sequence Generation
Several commercial CAD softwares [28] employ heuristic feature extraction to generate command sequences from B-Rep models. Fusion 360 Gallery [11] is a representative deep learning-based approach that first extracts features to obtain continuous geometric parameters for basic operations (e.g., 2D sketching, extrusion, etc.). It then uses imitation learning to generate confidence scores of the command sequence from the input B-Rep model. Finally, a confidence-based local search algorithm, combined with geometric similarity comparison, optimizes model accuracy in a simulation environment. However, this method only supports face-extrusion operations, and the local search algorithm heavily relies on the accuracy of the confidence scores from supervised learning.
@@ 3 Overview and Preliminary
Figure 2: RLCAD employs Graph Transformer Networks (GTN) to extract features from the current B-Rep and target B-Rep, obtaining feature vectors $g_{c}$ and $g_{t}$, respectively. These feature vectors $g_{c}$ and $g_{t}$ are then fused using self-attention. For the action sequence, a $L$-layer GTrXL is utilized to extract sequential features, and the last valid action $h_{\text{last\_valid}}$ is taken as the final feature of the sequence. Cross-attention is applied to fuse the features of $\tilde{G}_{\text{stack}}$ and $h_{\text{last\_valid}}$, followed by a MLP for feature transformation. The transformed features are then used as inputs to the Actor and Critic networks.
@@ 3.1 Our Approach
The overall training pipeline is depicted in Figure 2. The whole network can be divided into two parts: feature extraction network and Actor-Critic network. Inspired by advancements in natural language processing (NLP), we consider incorporating the action sequence as part of the state and fusing it with B-rep features using cross-attention. This allows us to measure the gap between the current action sequence and the target geometry. Specifically, for the current B-rep and target B-rep, we use Graph Transformer Networks (GTN) to extract features from the face adjacency graphs, obtaining $g_{c}$ and $g_{t}$. These features are stacked to form $G_{\text{stack}}$, and self-attention is applied to fuse $g_{c}$ and $g_{t}$, resulting in $\tilde{G}_{\text{stack}}$. For the action sequence, we use an $L$-layer GTrXL network to extract features. And then we take the last valid action $h_{\text{last\_valid}}$ as the feature of the action sequence. Cross-attention is then applied to fuse $h_{\text{last\_valid}}$ and $\tilde{G}_{\text{stack}}$. Finally, the fused feature is served as input to the Actor and Critic networks, predicting the probability distribution of actions and the cumulative reward of the current state, respectively. The action is sampled from the probability distribution and interacts with RLCADGym to obtain the reward and the next state. Through multiple rounds of interaction and network parameters update, the optimal action sequence is predicted.
@@ 3.2 Preliminary
We formulate the CAD reconstruction task from B-Rep to command sequence as a Markov Decision Process (MDP). The key components of this formulation are defined as follows: States $(s)\in\mathcal{S}$ contains the current and target B-Rep geometry, we use a face-adjacency graph [39] to represent the B-Rep geometry. Action $(a)\in\mathcal{A}$ represents a modeling operation that allows the simulator to modify the current geometry with extrusion, revolution and Boolean operations. State transition kernel $T$ allows the simulator to apply the modeling operation to update the current geometry. Reward $(r)\in\mathcal{R}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ allows the simulator to calculate the difference between the current and target states, and reward discount $\gamma\in[0,1)$, determining the importance of future rewards.
@@ 4 CAD Training Gym
We developed a CAD training gym based on the Parasolid geometric engine. By encapsulating the API of CAD engine, we design a set of high-level development interfaces for command-sequence-based modeling. Each interface corresponds to an extrusion or revolution operation combined with a Boolean operation (newbody, intersection, union, or subtraction) that incrementally modifies the geometry. The gym also supports parallel sampling across multiple scenarios, improving RL training efficiency.
@@ 5.1 Proximal Policy Optimization
$\widehat{A}_{t}=\sum_{k=0}^{|\tau|-t}(\gamma\lambda)^{k}\delta_{t+k}.$
@@ 5.3 Action Encoding
In CAD modeling tasks, efficient action representation is crucial for the performance of RL algorithms. We propose an index mapping method that directly maps CAD operations to discrete action space, effectively reducing the dimensionality of the state-action space. Each CAD operation is defined by four key components: $a=(f_{s},f_{e},o_{t},a_{t})$, where $o_{t}$ represents the operation type (including Newbody, intersection, union, and subtraction), $a_{t}$ indicates the action type (extrude and revolve), for extrude operation $f_{s}$ denotes the starting face ID, $f_{e}$ denotes the ending face ID; for revolve operation $f_{s}$ equals $f_{e}$.
@@ 5.4 Policy Network
Our policy network consists of feature extraction modules and Critic-Actor network. As shown in Figure  2, the network processes three inputs: B-Reps of the target and current models, and historical action sequences. First, we extract B-Rep modeling of the target and current models, encoding each B-Rep graph through an independent GTN with two TransformerConv layers  [43] to produce 256-dimensional embeddings $g_{t}$ (target) and $g_{c}$ (current). These embeddings are stacked into $G_{stack}\in\mathbb{R}^{2\times 256}$, which undergoes 8-head self-attention alignment, where $G_{stack}$ serves as both Query and Key/Value inputs. This generates the refined tensor $\tilde{G}_{stack}$. We then split $\tilde{G}_{stack}$ along the stacking dimension to obtain enhanced features $\tilde{g}_{t}$ and $\tilde{g}_{c}$, finally concatenating them into a fused geometric representation $G_{cat}\in\mathbb{R}^{512}$.
@@
Cross-modal fusion aligns geometric and action features through attention mechanisms: $h_{\text{last\_valid}}^{(L)}$ serves as Query while stacked geometric features $\tilde{G}_{stack}$ act as Key/Value, producing action-conditioned context $F_{a}\in\mathbb{R}^{256}$. The fused features $F_{a}$ and $G_{cat}$ are concatenated into a 768-dimensional vector, which is projected to 2048-dimensional through MLP.
@@ 5.5 Reward Shaping
However, observations reveal that relying solely on IoU led to suboptimal policies. The agent prioritized maximizing coarse volumetric overlap while neglecting fine geometric details. To address this limitation, we introduce two complementary metrics: Minimum Matching Distance (MMD) [45] and Normal Consistency (NC) [46], which explicitly enforce local geometric fidelity and surface quality.
@@
$\mathrm{R}_{MMD}=-\frac{1}{2}\left(\text{MMD-CD}(\mathcal{G},\mathcal{S})+\text{MMD-EMD}(\mathcal{G},\mathcal{S})\right),$
@@
$\mathrm{R}=\alpha\cdot\text{IoU}(\mathcal{G},\mathcal{S})+\beta\cdot\text{R}_{MMD}+\gamma\cdot\text{NC}(\mathcal{G},\mathcal{S}),$
where the weights $\alpha=0.2$, $\beta=0.5$, and $\gamma=0.3$ were determined through iterative experiments to balance global shape alignment $\alpha$, local geometric fidelity $\beta$, and surface quality $\gamma$.
This hierarchical design aligns with the progressive nature of CAD modeling: IoU guides coarse volumetric exploration in the early stages, while MMD and NC refine geometric details as the model converges.
@@ 6 Experiments and Comparison
We evaluate our CAD command sequence generation algorithm on a workstation equipped with 2 × 16-core AMD Ryzen 9 9950X CPUs and an Nvidia RTX 3090 GPU (24GB memory).
@@ 6.1 Evaluation Metrics
To assess the quality of the generated CAD models, we first convert the B-Rep models into point clouds. Let $\mathcal{S}$ denote the reference models and $\mathcal{G}$ represent the generated models. We employ IoU, COV, MMD-CD, JSD and NC as metrics to evaluate the quality of our generated CAD models [45]. IoU, MMD-CD, and NC have been introduced in Section 5.5.
@@
$\text{COV}(\mathcal{G},\mathcal{S})=\frac{|\{\arg\min_{\mathcal{Y}\in\mathcal{S}}d_{CD}(\mathcal{X},\mathcal{Y})|\mathcal{X}\in\mathcal{G}\}|}{|\mathcal{S}|}$ (1)
Jensen-Shannon Divergence (JSD) measures the dissimilarity between two point clouds from the perspective of voxel distribution.
$\text{JSD}(P_{\mathcal{G}},P_{\mathcal{S}})=\frac{1}{2}D(P_{\mathcal{S}}||M)+\frac{1}{2}D(P_{\mathcal{G}}||M),$ (2)

### v2

@@ Abstract
In this paper, we present a reinforcement learning (RL) training gym specifically designed for CAD model generation, along with an RL-based algorithm that generates command sequences from boundary representation (B-Rep) geometry within this training gym. Given an input B-Rep, the policy network of the RL algorithm first outputs an action. This action, together with previously generated actions, is processed within the gym to produce the corresponding CAD geometry, which is then fed back into the policy network. Rewards, computed by the difference between the generated and target geometries within the gym, are used to update the RL network. Our method supports operations beyond sketches, Boolean, and extrusion, including revolution operations. With this training gym, we achieve state-of-the-art (SOTA) quality in generating command sequences from B-Rep geometries.
@@ 1 Introduction
Figure 1: Demonstration of various CAD models generated by our novel CAD model generation algorithm. Including revolution operation enables the generation of more complex geometric shapes, such as the battery, nuts, flanges, etc.
In our daily lives and industrial production, objects ranging from small items like cups to large structures like airplanes are designed using a series of 2D and 3D modeling operations in computer-aided design (CAD) software. The process of recording these parametric operations results in a command sequence that encapsulates domain expert knowledge by precisely defining a sequence of modeling steps to generate CAD geometries. However, obtaining the command sequence directly is often not feasible. Instead, we should generate the corresponding command sequence using other modalities, such as boundary representation, meshes, or point clouds.
@@
Traditional CAD geometry generation algorithms rely on CAD geometric engines, which provide essential functionalities such as validity checks and constraint solving – capabilities that supervised learning alone cannot achieve. Meanwhile, SOTA large language models such as DeepSeek R1 [19] and OpenAI o1 [20] employ reinforcement learning (RL) techniques to enhance text generation quality and model reasoning capabilities. In bionic robotics, RL-based simulation training has also yielded impressive results in robot motion control [21] and manipulation [22].
Inspired by these advancements, we present RLCAD, a CAD RL training gym based on CAD geometric engine. In this framework, the policy network of the RL framework generates a command sequence, which is executed in the gym to produce the corresponding CAD geometry. This geometry is then fed back into the policy network to calculate rewards and update the network.
Based on this training gym, we propose an RL-based CAD geometry generation algorithm that converts B-Rep models into command sequences. Our approach supports operations beyond sketches, Boolean, and extrusion, including revolution operations. As shown in Figure 1, including revolution operation enables the generation of more complex geometric shapes, such as the battery, nuts, chess pieces, etc. We summarize our contributions as follows:
@@
    We present a CAD RL training gym bridging RL algorithms with CAD geometric engines. We provide a high-level Python interface wrapper and support parallel environment sampling, enabling efficient interaction with multiple CAD environments. In addition, we integrate the RL algorithm library Stable Baseline3 [23] and the OpenAI Gymnasium interface [24] with our training gym, making it easier for developers to implement and test RL algorithms.
@@
    We present an RL-based CAD model generation algorithm that converts B-Rep models into command sequences within our training gym. Our approach features a multi-modal policy network that integrates multiple common CAD operations with cross-modal feature extractors. To further enhance generation quality, we introduce a set of hybrid reward functions that jointly guide the learning process.
@@ 2.1 CAD Datasets
Current parametric CAD models typically are stored as B-Rep [25] or command sequence [26]. B-Rep includes the analytical representation of the surface and curves of a 3D shape, while command sequences record the historical construction process of the shape. B-Rep models consist of trimmed parametric surfaces along with topological information describing their adjacency relationships.
@@ 2.2 CAD Model Generation
BrepGen [7] combines vector variational autoencoders and latent space diffusion models for the autoregressive generation of B-Rep models. CADSIGNet [8] employs layer-wise cross-attention between point clouds and command sequences to learn implicit representations of CAD models, enabling end-to-end point cloud to command sequence generation. HNC-CAD [14], SkexGen [15], and DeepCAD [16] quantize and encode CAD command sequences, generating corresponding latent space features and training through an encoder-decoder autoregressive model. SfmCAD [13] uses voxels as input and proposes differentiable sketch and sweeping path modeling operations to reconstruct the CAD model in a self-supervised manner.
@@ 2.3 B-Rep to Command Sequence Generation
Several commercial CAD softwares [28] employ heuristic feature extraction to generate command sequences from B-Rep models. However, the 2D sketches extracted using this method are often structurally complex and difficult to edit further. Fusion 360 Gallery [11] is a representative deep learning-based approach that first extracts features to obtain continuous geometric parameters for basic operations (e.g., 2D sketching, extrusion, etc.). It then uses imitation learning to generate confidence scores of the command sequence from the input B-Rep model. Finally, a confidence-based local search algorithm, combined with geometric similarity comparison, optimizes model accuracy in a simulation environment. However, this method only supports face-extrusion operations, and the local search algorithm heavily relies on the accuracy of the confidence scores from supervised learning.
@@ 2.4 RL Gym
Figure 2: The training pipeline is composed of two stages. In the first stage, a contrastive learning approach is employed to pre-train the UV-Net network, aiming to derive an encoder model that can effectively characterize the B-Rep of CAD model. During the second stage, a reinforcement learning approach is employed to generate the command sequence. We first utilize the tunable UV-Net model to extract the B-Rep embedding of the CAD model, which is then integrated with the feature vector of the historical modeling action sequence. Subsequently, the Actor-Critic network predicts the action distribution and value. The predicted action is transmitted to RLCADGym for execution, yielding the next-stage observation. The neural reward and geometric reward are designed to update the policy network.
@@ 3.1 Our Approach
The overall training pipeline is depicted in Figure 2. In the first stage, we adopt a contrastive learning approach to pre-train UV-Net [39] on a large-scale dataset of unlabeled CAD models, aiming to derive an embedding that can effectively characterize the high-dimensional B-Rep feature of CAD model. Given a B-Rep model represented as a face adjacency graph, UV-Net processes it through a combination of surface convolution and curve convolution operations, followed by a Graph Neural Network (GNN) to extract a compact and discriminative embedding. To enhance the robustness of the learned representations, we apply UV-Net’s data augmentation strategies, including connected patch extraction and stochastic node/edge dropping.
In the second training stage, the whole network can be divided into two parts: the feature extraction network and the Actor-Critic network. Inspired by advancements in natural language processing (NLP), we consider incorporating the action sequence as part of the state and fusing it with B-Rep features using cross-attention. This allows us to measure the gap between the current action sequence and the target geometry. Specifically, for the current B-Rep and target B-Rep, we use the pre-trained UV-Net to extract features from the face adjacency graphs, obtaining $g_{c}$ and $g_{t}$. These features are stacked to form $G_{\text{stack}}$, and self-attention is applied to fuse $g_{c}$ and $g_{t}$, resulting in $\tilde{G}_{\text{stack}}$. For the action sequence, we use an $L$-layer GTrXL network to extract features. And then we take the last valid action $h_{\text{last\_valid}}$ as the feature of the action sequence. Cross-attention is then applied to fuse $h_{\text{last\_valid}}$ and $\tilde{G}_{\text{stack}}$. Finally, the fused feature is served as input to the Actor and Critic networks, predicting the probability distribution of actions and the cumulative reward of the current state, respectively. The action is sampled from the probability distribution and interacts with RLCADGym to obtain the reward and the next state. Through multiple rounds of interaction and network parameters update, the optimal action sequence is predicted.
@@ 3.2 Preliminary
We formulate the CAD reconstruction task from B-Rep to command sequence as a Markov Decision Process (MDP). The key components of this formulation are defined as follows: States $(s)\in\mathcal{S}$ contains the current and target B-Rep geometry, we use a face-adjacency graph [39] to represent the B-Rep geometry. Action $(a)\in\mathcal{A}$ represents a modeling operation that allows the simulator to modify the current geometry with extrusion, revolution, and Boolean operations. State transition kernel $T$ allows the simulator to apply the modeling operation to update the current geometry. Reward $(r)\in\mathcal{R}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ allows the simulator to calculate the difference between the current and target states, and reward discount $\gamma\in[0,1)$, determining the importance of future rewards.
@@ 4 CAD Training Gym
We develop a CAD training gym based on the Parasolid geometric engine. By encapsulating the API of the CAD engine, we design a set of high-level development interfaces for command-sequence-based modeling. Each interface corresponds to an extrusion or revolution operation combined with a Boolean operation (newbody, intersection, union, or subtraction) that incrementally modifies the geometry. The gym also supports parallel sampling across multiple scenarios, improving RL training efficiency.
@@ 5.1 Proximal Policy Optimization
$\widehat{A}_{t}=\sum_{k=0}^{|\tau|-t}(\gamma\lambda)^{k}\delta_{t+k},$
@@ 5.3 Action Encoding
In CAD modeling tasks, efficient action representation is crucial for the performance of RL algorithms. We propose an index mapping method that directly maps CAD operations to a discrete action space, effectively reducing the dimensionality of the state-action space. Each CAD operation is defined by four key components: $a=(f_{s},f_{e},o_{t},a_{t})$, where $o_{t}$ represents the operation type (including newbody, intersection, union, and subtraction), $a_{t}$ indicates the action type (extrude and revolve), for extrude operation $f_{s}$ denotes the starting face ID, $f_{e}$ denotes the ending face ID; for revolve operation $f_{s}$ equals $f_{e}$.
@@ 5.4 Policy Network
Our policy network consists of feature extraction modules and a Critic-Actor network. As shown in Figure  2, the network processes three inputs: B-Reps of the target and current models, and historical action sequences. First, we extract B-Rep modeling of the target and current models, encoding each B-Rep graph through an independent UV-Net to produce 256-dimensional embeddings $g_{t}$ (target) and $g_{c}$ (current). These embeddings are stacked into $G_{stack}\in\mathbb{R}^{2\times 256}$, which undergoes 8-head self-attention alignment, where $G_{stack}$ serves as both Query and Key/Value inputs. This generates the refined tensor $\tilde{G}_{stack}$. We then split $\tilde{G}_{stack}$ along the stacking dimension to obtain enhanced features $\tilde{g}_{t}$ and $\tilde{g}_{c}$, finally concatenating them into a fused geometric representation $G_{cat}\in\mathbb{R}^{512}$.
@@
Cross-modal fusion aligns geometric and action features through attention mechanisms: $h_{\text{last\_valid}}^{(L)}$ serves as Query while stacked geometric features $\tilde{G}_{stack}$ act as Key/Value, producing action-conditioned context $F_{a}\in\mathbb{R}^{256}$. The fused features $F_{a}$ and $G_{cat}$ are concatenated into a 768-dimensional vector, which is projected to 2048 dimensions through an MLP.
@@ 5.5 Reward Shaping
However, observations reveal that relying solely on IoU led to suboptimal policies. The agent prioritized maximizing coarse volumetric overlap while neglecting fine geometric details. To address this limitation, we introduce three complementary rewards: Minimum Matching Distance (MMD) [44], Normal Consistency (NC) [45] and Neural Reward (NR) which explicitly enforce local geometric fidelity, surface quality and high-dimensional feature similarity.
@@
$\mathrm{MMD}=-\frac{1}{2}\left(\text{MMD-CD}(\mathcal{G},\mathcal{S})+\text{MMD-EMD}(\mathcal{G},\mathcal{S})\right),$
@@
Neural Reward (NR) addresses the limitations of conventional geometric rewards, which have limited sensitivity to subtle CAD features. NR leverages UV-Net’s B-Rep embeddings: by reformulating rewards as cosine similarities between these high-dimensional features, it enables better feature-aware optimization. It is formulated as:
$\mathrm{NR}(\mathcal{G},\mathcal{S})=\cos\left(f_{UV-Net}\left(\mathcal{G}\right),f_{UV-Net}\left(\mathcal{S}\right)\right).$
@@
$\mathrm{R}$ $=~{}\alpha\cdot\text{IoU}(\mathcal{G},\mathcal{S})+\beta\cdot\text{MMD}+\gamma\cdot\text{NC}(\mathcal{G},\mathcal{S})$
$+\delta\cdot\text{NR}(\mathcal{G},\mathcal{S}).$
where the weights $\alpha=0.3$, $\beta=0.2$, $\gamma=0.2$, and $\delta=0.3$ are determined empirically to balance global shape alignment (IoU), local geometric fidelity (MMD), surface quality (NC), and semantic similarity (NR).
@@ 6 Experiments and Comparison
We evaluate our CAD command sequence generation algorithm on a workstation equipped with an NVIDIA L20 GPU (48GB memory).
Datasets
We construct a CAD dataset containing extrusion and revolution operations for training and evaluation, which includes $20k$ CAD models from the ABC dataset  [27] after online parsing and deduplication. The dataset is stratified by the number of faces to measure model complexity, comprising $14k$ simple models (with fewer than 10 faces), $4k$ medium models (with 10 to 20 faces), and $2k$ complex models (with more than 20 faces). Using stratified sampling, we divide the dataset into a training set of $19k$ models and a testing set of $1k$ models, where complex models with over 20 faces account for 39.8% of the test dataset.
To further enhance the pre-training performance, we employ the DeepCAD dataset ($170k$ samples) in conjunction with our training dataset ($19k$ samples), forming a comprehensive pre-training corpus of $189k$ CAD models. We analyze the geometric complexity of the pre-training dataset in terms of face count per model: 33% contain fewer than 10 faces, 32% have 10–20, 12% have 20–30, 6% have 30–40, and 17% contain more than 40 faces.
@@ 6.1 Evaluation Metrics
To assess the quality of the generated CAD models, we first convert the B-Rep models into point clouds. Let $\mathcal{S}$ denote the reference models and $\mathcal{G}$ represent the generated models. We employ IoU, COV, MMD-CD, JSD, and NC as metrics to evaluate the quality of our generated CAD models [44]. IoU, MMD-CD, and NC have been introduced in Section 5.5.
@@
$\text{COV}(\mathcal{G},\mathcal{S})=\frac{|\{\arg\min_{\mathcal{Y}\in\mathcal{S}}d_{CD}(\mathcal{X},\mathcal{Y})|\mathcal{X}\in\mathcal{G}\}|}{|\mathcal{S}|}.$
Jensen-Shannon Divergence (JSD) measures the dissimilarity between two point clouds from the perspective of voxel distribution:
$\text{JSD}(P_{\mathcal{G}},P_{\mathcal{S}})=\frac{1}{2}D(P_{\mathcal{S}}||M)+\frac{1}{2}D(P_{\mathcal{G}}||M),$
@@
6.2 Training Efficiency Comparison
Fusion 360 provides a server-side plugin called Fusion 360 Gym [11], which sequentially receives CAD modeling commands sent from external programs and returns the results to those programs via Hypertext Transfer Protocol (HTTP). Based on this plugin, we connect our reinforcement learning algorithm with Fusion 360 to enable closed-loop training.
We compared our novel training gym with Fusion 360 Gym over 10,000 training steps. Using parallel computation (1, 8, and 16 environments), we evaluated both memory consumption and execution time per step. Table 1 shows that Fusion 360 Gym consumes substantially more memory. The frequent use of the revert operation in Fusion 360 introduces a risk of memory overflow. In contrast, our gym employs a stable mark-and-revert mechanism that maintains steady memory usage without significant fluctuations. In addition, by replacing Fusion 360’s network communication mechanism with multi-process parallelism and shared memory architecture, we achieve a significant reduction in data transmission latency between the reinforcement learning algorithm and the gym. We get up to $39X$ speedup compared with Fusion 360 Gym, as shown in Table 2.
Table 1: Comparison of memory usage between Fusion 360 and our gym (unit: MB). It shows that Fusion 360 Gym consumes substantially more memory, while our training gym maintains steady memory usage.
Fusion 1-env 8-env 16-env
Initial Memory 1203.6 9.3 10.4 9.5
Final Memory 6362.7 10.2 11.8 10.3
Table 2: Comparison of execution time per step (unit: ms). It shows that our method get up to $39X$ speedup compared with Fusion 360 Gym.
Fusion 1-env 8-env 16-env
Execution Time 1674 198 52.5 42.18
Speedup - 8.45 31.89 39.69